# Multiscale Quantile Regression with Local Error Control

### v2

@@ 5.1.1 Changes in median
  • MUSCLE is among the best ones in all scenarios, and in particular, outperforms the others by a large margin in scenarios (E4) and (E5). The competitive performance of MUSCLE aligns with its statistical guarantees (Section 3). As a speedup variant of MUSCLE, MUSCLE-S has nearly the same performance as MUSCLE, and tends to report slightly more change points.
Regarding computation time, PELT and R-FPOP are the fastest, followed closely by SMUCE, WBS, NOT-HT, MUSCLE-S, MQS and FDRSeg. MUSCLE is noticeably slower, and RNSP and KSD are the slowest. In all scenarios, MUSCLE-S is roughly ten times faster than MUSCLE, and takes about 1 s for 2000 samples. See Table 2 in Appendix B.
Figure 3: Recovery of the underlying signal in (E1). In each panel, the overall performance over 200 repetitions is summarised as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true number of change points is $K=2$ (marked by a red dashed line).
Figure 4: Recovery of the blocks signal in (E2). In each panel, the overall performance over 200 repetitions is summarised as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true numbers of change points in median and in distribution are $K=11$ (marked by a red dashed line) and $K=14$ (marked by a blue dashed line), respectively.
Figure 5: Recovery of the blocks signal in (E5). In each panel, the overall performance over 200 repetitions is summarised as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true numbers of change points in median and in distribution are $K=11$ (marked by a red dashed line) and $K=14$ (marked by a blue dashed line), respectively.
@@ 5.1.2 Changes in distribution
[…]
$\mathcal{C}_{k}^{m}\coloneqq\biggl{\{}g=\sum_{j=0}^{k}\theta_{j}\mathbbm{1}_{{I}_{j}}:\theta_{j}=(\theta_{j1},\ldots,\theta_{jm}),\bigsqcup_{j=0}^{k}I_{j}=[0,1),\\ \text{ and }T_{\accentset{\circ}{I}_{j}}(Z,\theta_{jr})\leq q_{\alpha/m}(\lvert\accentset{\circ}{I}_{j}\rvert),\,j=0,\dots,k,\,r=1,\ldots,m\biggr{\}}.$
@@
We demonstrate the performance of M-MUSCLE in scenario (E2) with $t_{3}$-noises, see Figure 6. We select $\beta=(\beta_{1},\beta_{2},\beta_{3})=(0.25,0.5,0.75)$ and set $\alpha=0.3$. M-MUSCLE estimates quite accurately the number of change points as well as their locations. In comparison with other nonparametric methods that aim at distributional changes, KSD often underestimates the number of change points, while ED-PELT largely overestimates the number of change points (see Figure 4, the bottom right panel).
Figure 6: Recovery of the blocks signal in (E2) by M-MUSCLE ($\alpha=0.3$, bottom). The estimated upper and lower quartiles are in orange lines and the estimated median in red line. The histogram on the right summarises its performance in estimating the number of change points over 200 repetitions, where the vertical dashed red line indicates the true number of change points. On the top, the true signal (black line) and the data (grey line) are plotted. Three change points of variances are highlighted by vertical dashed blue lines.
@@ 5.2.1 Well-log dataset
The well-log dataset (Ruanaidh and Fitzgerald, , 2012) contains 4050 measurements of the nuclear magnetic response of underground rocks. Collected by inserting a probe into a borehole, the true signal exhibits approximate piecewise constancy, with each segment corresponding to a distinct rock type characterised by constant physical properties. Change points in the signal reveal changes between rock types, holding crucial importance in the context of oil drilling. We refer to Fearnhead, (2006), Wyse et al., (2011) and Fearnhead and Rigaill, (2019) for further details.
@@
The raw data exhibits several short but sharp oscillations, with some of these fluctuations significantly smaller in amplitude compared to others. Given these distinct patterns, a Gaussian noise assumption might not be justified, which is supported by Shapiro–Wilk’s test, yielding a $p$-value of $2.2\times 10^{-16}$. Consequently, many studies analysing well-log data rely on a pre-cleaned dataset, wherein outliers are manually removed in advance. In general, the detection and removal of outliers are challenging tasks due to the scarcity of labelled data, stemming from the infrequent occurrence of outlier instances. Typical outlier detection methods identify data points that deviate significantly from the rest. However, proposing a universal measure of deviation applicable to all datasets and scenarios is hard or even impossible (cf. Boukerche et al., , 2020).
@@
The robustness of MUSCLE eliminates the need for a preliminary pre-cleaning step. We employ MUSCLE to estimate the median ($\beta=1/2$) of well-log signal with different choices of the tuning parameter, namely, $\alpha=0.1,0.3$ and $0.5$, and by comparison we also consider other robust or nonparametric segmentation methods, including KSD, MQS, RNSP, NOT-HT, R-FPOP and ED-PELT, see Figure 7. As shown in the top panel, the raw data has five pronounced flickering events (i.e., events on tiny temporal scales). The estimates provided by KSD, MQS, RNSP and MUSCLE with $\alpha=0.1$ or $0.3$ are very similar, and none of them detect any flickering events. Despite this, they all capture the main shape of the well-log data. We set $\alpha=0.5$ for MQS and RNSP in order to improve their detection power. By contrast, NOT-HT identifies two and R-FPOP detects three flickering events. However, R-FPOP seems to overfit the data around the last flickering (see Figure 11 in Appendix B). ED-PELT reports the highest number of change points. However, as it constantly overestimates $K$ in simulation studies, its estimation may not be trustworthy. With $\alpha=0.5$, MUSCLE successfully recovers all of the five flickering events. In general, it is challenging to verify whether these flickering events (or some of them) are genuine change points or outliers causing by mistakes in measurement. Nevertheless, note that MUSCLE with various choices of $\alpha$ reveals a nested sets of detected features in the well-log data. As a larger $\alpha$ corresponds a lower statistical confidence, such a nested sets of detected features shed light on how likely these features are genuine.
@@ 5.2.2 Ion channels recordings: Gramicidin D
with $\omega$ the sampling rate, and $\rho$ the convolution kernel (corresponding to the low-pass filter). The underlying signal $f$ represents the conductance profile of ion channels, which is a piecewise constant function with change points caused by gating events. The noises $\varepsilon_{i}$ are centred, heterogeneous, and typically non-Gaussian (Venkataramanan et al., , 1998), and they are dependent due to the low-pass filter.
We showcase an extension of MUSCLE to the dependent noises in (17). To take into account the dependence structure, we need to adjust the local quantiles $q_{\alpha}$ in (4). It is generally difficult to determine the dependence of $\mathbbm{1}_{\{{\varepsilon}_{i}\leq 0\}}$, even though the dependence of noises $\varepsilon_{i}$ arises from the known low-pass filter, unless specific assumptions on the noise distribution function are made. Note however that, as a crucial building block of MUSCLE, the multiscale test statistic converges weakly to an accessible distribution in form of supremum of Gaussian random variables under mild conditions (see e.g. Frick et al., , 2014 and Dette et al., , 2020). Inspired by this, we redefine the local quantiles $q_{\alpha}$ in (4) by treating noises $\varepsilon_{i}$ as dependent Gaussian distributed, with the dependence structure derived from the convolution kernel $\rho$. We compute such local quantiles by Monte–Carlo simulations. Another challenge posed by model (17) is that the convoluted signal $\rho*f$ is no longer piecewise constant. But the influence of convolution is only locally active, as the convolution kernel $\rho$ has a compact support. As a consequence, $\rho*f$ is still constant on $[s+\kappa,t)$, with $\kappa$ the kernel length (i.e. the size of the support of $\rho$), if $f$ is constant on $[s,t)$. Thus, we also modify the multiscale statistic $T_{I}$ in (3) by restricting to intervals $J$ such that $J\subseteq I$ and $J-\kappa\subseteq I$. Based on these two adjustments, we obtain an extension of MUSCLE for dependent errors, denoted by D-MUSCLE.
@@ 6 Discussion
In this paper, we introduce a novel change point segmentation method MUSCLE that is robust and meanwhile has a high detection power. It combines the ideas of multiscale testing, quantile segmentation and variational estimation. As a distinctive feature, MUSCLE controls the local errors associated with individual segments. This local error control, as opposed to the commonly studied global error control, allows improved detection power, robustness to data windowing, as well as efficient computation. Under the quantile segmentation model with serial independence, we have established the consistency and the localisation error rates for MUSCLE. Remarkably, these results match (up to log factors) the minimax optimality results in Gaussian setups. Further, we obtain finite sample guarantees on its falsely detected change points in terms of FDR and OER. This provides a statistical interpretation on the unique tuning parameter $\alpha$ of MUSCLE in the sense that the expected proportion of falsely detected change points does not exceed $\alpha$. Moreover, we develop an efficient dynamic programming algorithm for MUSCLE, which leverages advanced wavelet tree data structures, modern pruning techniques, and sparse interval system proposals. We also introduce a speedup variant (i.e., MUSCLE-S) of MUSCLE, which scales linearly with sample size, particularly suitable for large scale datasets. The efficacy of MUSCLE(-S) is validated through comprehensive comparison studies using both simulated and real world data.
@@
7 Acknowledgements
This work was supported by DFG (German Reserach Foundation) CRC 1456 Mathematics of Experiment, by DFG FOR 5381 Mathematical Statistics in the Information Age, and by the DFG under Germany’s Excellence Strategy, project EXC 2067 Multiscale Bioimaging: from Molecular Machines to Networks of Excitable Cells (MBExC). The authors thank Manuel Fink and Claudia Steinem (Institute of Organic and Biomolecular Chemistry, University of Göttingen) for providing the ion channel data, and also Axel Munk and Robin Requadt for helpful discussions.
Appendix A Proof of Theorems
@@ A.1 Technical results
$A_{k}(x)$ $\coloneqq\frac{x}{k}\log\left(\frac{x}{k}\cdot\frac{1}{\beta}\right)+\left(1-\frac{x}{k}\right)\log\left(\left(1-\frac{x}{k}\right)\cdot\frac{1}{1-\beta}\right),$
@@
Define a series of random times:
$T_{0}$ $\coloneqq 0,$
$T_{s}$ $\coloneqq\max\left\{l\,\bigg{|}\,\underset{T_{s-1}<i<j\leq l}{\max}B_{l-T_{s-1},i,j}-q_{\alpha}(l-T_{s-1})\leq 0\right\},$
@@
Let $S$ be in (18) and $\alpha\in(0,1)$. Then for any $s\in\mathbb{N}$,
$\mathbb{P}(S\geq s)\leq\alpha^{s}.$
Moreover, we have
$\mathbb{E}(S)\leq\frac{\alpha}{1-\alpha}.$
@@
$\mathbb{P}(T_{s}-T_{s-1}\geq n\,|\,T_{s-1}=t_{s-1},\dots,T_{1}=t_{1})$
$=$ $\,\mathbb{P}\left(\underset{t_{s-1}<i<j\leq l}{\max}B_{l-k,i,j}-q_{\alpha}(l-t_{s-1})\leq 0\text{ for some }l\geq n+t_{s-1}\right)$
$\geq$  $\,1-\alpha.$ (19)
Note that
$\mathbb{P}(S\geq s)=\mathbb{P}(T_{s}<n)=\mathbb{P}\left(\sum_{i=1}^{s}T_{s}<n\right)\leq\mathbb{P}(T_{i}-T_{i-1}<n\text{ for all }i=1,\dots,s).$
We then show by introduction over $s$ that for any $n\geq 1$, $\mathbb{P}(T_{i}-T_{i-1}<n\text{ for all }i=1,\dots,s)\leq\alpha^{s}$. For $s=1$, by definitions of $q_{\alpha}(m)$ in (4) we have $\mathbb{P}(T_{1}-T_{0}<n)=\mathbb{P}(T_{1}<n)=1-\mathbb{P}(T_{1}\geq n)\leq 1-(1-\alpha)=\alpha$. Now assume that the statement holds for $s-1$, i.e., $\mathbb{P}(T_{i}-T_{i-1}<n\text{ for all }i=1,\dots,s-1)\leq\alpha^{s-1}$. Then
@@
$=$ $\,\sum_{t_{i}-t_{i-1}<n,\,i=1,\dots,s-1}\mathbb{P}(T_{s}-T_{s}-1<n,\,T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\,\dots,\,T_{2}-T_{1}=t_{2}-t_{1},\,T_{1}=t_{1})$
$=$ $\,\sum_{t_{i}-t_{i-1}<n,\,i=1,\dots,s-1}\mathbb{P}(T_{s}-T_{s}-1<n\,|\,T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\,\dots,\,T_{2}-T_{1}=t_{2}-t_{1},\,T_{1}=t_{1})$
×$\qquad\qquad\qquad\qquad\times\mathbb{P}(T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\,\dots,\,T_{2}-T_{1}=t_{2}-t_{1},\,T_{1}=t_{1})$
$=$ $\,\sum_{t_{i}-t_{i-1}<n,\,i=1,\dots,s-1}\big{(}1-\mathbb{P}(T_{s}-T_{s}-1\geq n\,|\,T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\,\dots,\,T_{2}-T_{1}=t_{2}-t_{1},\,T_{1}=t_{1})\big{)}$
×$\qquad\qquad\qquad\qquad\times\mathbb{P}(T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\,\dots,\,T_{2}-T_{1}=t_{2}-t_{1},\,T_{1}=t_{1})$
$\leq$ $\,\sum_{t_{i}-t_{i-1}<n,\,i=1,\dots,s-1}\alpha\cdot\mathbb{P}(T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\,\dots,\,T_{2}-T_{1}=t_{2}-t_{1},\,T_{1}=t_{1})$
@@
Let $\beta\in(0,1)$ and suppose that $Z_{1},\dots,Z_{n}$ have the common $\beta$-quantile, i.e., ${\mathbb{P}(Z_{i}\leq f(x_{i}))=\beta}$ with ${f}\equiv\theta$ for some $\theta\in\mathbb{R}$. Then for any $k\geq 1$, the estimated number of change point $\widehat{K}$ by MUSCLE satisfies
$\mathbb{P}\left(\widehat{K}\geq k\right)$ $\leq\alpha^{k},$ (20)
$\text{and}\quad\mathbb{E}\left(\widehat{K}\right)$ $\leq\frac{\alpha}{1-\alpha}.$ (21)
@@
With loss of generality, we assume that ${f}=0$. Suppose that there exists an estimator $\widetilde{f}=\sum_{k=0}^{\widetilde{K}}\theta_{k}I_{k}$ with $\theta_{k-1}\neq\theta_{k}$ for $k=1,\dots,K$, which has $\widetilde{K}$ change points and satisfies the multiscale side-constrains of $C_{\tilde{K}}$ in (5), then $\widetilde{f}$ has exactly $\widetilde{K}$ false discoveries. Since the MUSCLE minimises the number of estimated change points fulfilling the multiscale side-constrains, we have $\widehat{K}\leq\widetilde{K}$. We now construct such an estimator $\widetilde{f}$ as follows.
@@
Set $T_{0}=0$. We fit the observations $\{Z_{i}\,|\,i\geq 1\}$ using the constant value 00 up to $T_{1}$, defined as the largest index $j$ such that 00 satisfies the multiscale constraint on the interval $[0,j/n)$. We then fit the subsequent observations $\{Z_{T_{1}+i}\,|\,i\geq 1\}$ using 00 until $T_{2}$ , the largest index $j$ such that 00 satisfies the multiscale constraint on $[T_{1}/n,j/n)$. Repeating this procedure until all observations are fitted yields a series of segments. For ensuring the resulting estimator has exactly $\widetilde{K}$ change points, we slightly perturb the segment values so that the multiscale constraints remain satisfied on each segment. More precisely, Let $W_{i}=\mathbbm{1}_{\{Z_{i}\leq 0\}}$ for $i=1,\dots,n$ and

### v3

@@ 5.1.1 Changes in median
  • MUSCLE is among the best ones in all scenarios, and in particular, in scenarios (E4) and (E5). The competitive performance of MUSCLE aligns with its statistical guarantees (Section 3). As a speedup variant of MUSCLE, MUSCLE-S has nearly the same performance as MUSCLE, and tends to report slightly more change points.
Regarding computation time (Table 2), PELT and R-FPOP are the fastest, followed closely by SMUCE, WBS, NOT-HT, MUSCLE-S, MQS and FDRSeg. MUSCLE is noticeably slower, and RNSP and KSD are the slowest. In all scenarios, MUSCLE-S is roughly ten times faster than MUSCLE, and takes about one second for 2000 samples. All runtimes are measured on a standard laptop (Intel i5 CPU, 1.40 GHz, four cores, 16 GB RAM).
Figure 3: Recovery of the underlying signal in (E1). In each panel, the overall performance over 200 repetitions is summarized as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true number of change points is $K=2$ (marked by a red dashed line).
Figure 4: Recovery of the blocks signal in (E2). In each panel, the overall performance over 200 repetitions is summarized as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true numbers of change points in median and in distribution are $K=11$ (marked by a red dashed line) and $K=14$ (marked by a blue dashed line), respectively.
Figure 5: Recovery of the blocks signal in (E5). In each panel, the overall performance over 200 repetitions is summarized as a boxplot and individual repetitions are jittered in dots with a low intensity. In the bottom left panel, the theoretical upper bound $\alpha=0.3$ on the FDR of MUSCLE is marked by a red dashed line. In the bottom right panel, the true numbers of change points in median and in distribution are $K=11$ (marked by a red dashed line) and $K=14$ (marked by a blue dashed line), respectively.
@@ 5.1.2 Changes in distribution
We now consider the detection of change points in distribution. If two random variables do not share the same distribution, at least one quantile of them must be different. Thus, a straightforward approach to detect change points in distribution is applying MUSCLE repeatedly to estimate various quantiles of observations, e.g., median ($\beta=0.5$), lower $(\beta=0.25)$ and upper ($\beta=0.75$) quartiles. However, for a change point in distribution, this straightforward approach may report several slightly different estimates when using different quantiles. To address this issue, we modify the multiscale tests on individual segments (see Section 2.1) by considering simultaneously multiple quantiles. Let $\beta=(\beta_{1},\ldots,\beta_{m})\in(0,1)^{m}$ be the considered levels of quantiles. For an interval $I\subseteq[0,\,1)$, we perform a multiscale test on whether $\{Z_{i}:x_{i}\equiv(i-1)/n\in I\}$ have the same $\beta_{r}$-quantile for every $r=1,\ldots,m$. Imitating the development in Section 2, we may arrive at the multiscale side-constraint
$\mathcal{C}_{k}^{m}\coloneqq\biggl\{g=\sum_{j=0}^{k}\theta_{j}\mathbbm{1}_{{I}_{j}}:\theta_{j}=(\theta_{j1},\ldots,\theta_{jm}),\bigsqcup_{j=0}^{k}I_{j}=[0,1),\\ \text{ and }T_{\mathring{I}_{j}}(Z,\theta_{jr})\leq q_{\alpha/m}(\lvert\mathring{I}_{j}\rvert),\,j=0,\dots,k,\,r=1,\ldots,m\biggr\}.$
@@
We demonstrate the performance of M-MUSCLE in scenario (E2) with $t_{3}$-noise, see Figure 6. We select $\beta=(\beta_{1},\beta_{2},\beta_{3})=(0.25,0.5,0.75)$ and set $\alpha=0.3$. M-MUSCLE estimates quite accurately the number of change points as well as their locations. In comparison with other nonparametric methods that aim at distributional changes, KSD often underestimates the number of change points, while ED-PELT largely overestimates the number of change points (see Figure 4, the bottom right panel).
Figure 6: Recovery of the blocks signal in (E2) by M-MUSCLE ($\alpha=0.3$, bottom). The estimated upper and lower quartiles are in orange lines and the estimated median in red line. The histogram on the right summarizes its performance in estimating the number of change points over 200 repetitions, where the vertical dashed red line indicates the true number of change points. On the top, the true signal (black line) and the data (grey line) are plotted. Three change points of variances are highlighted by vertical dashed blue lines.
@@ 5.2.1 Well-log dataset
The well-log dataset (Ruanaidh and Fitzgerald, , 2012) contains 4050 measurements of the nuclear magnetic response of underground rocks. Collected by inserting a probe into a borehole, the true signal exhibits approximate piecewise constancy, with each segment corresponding to a distinct rock type characterized by constant physical properties. Change points in the signal reveal changes between rock types, holding crucial importance in the context of oil drilling. We refer to Fearnhead, (2006), Wyse et al., (2011) and Fearnhead and Rigaill, (2019) for further details.
@@
The raw data exhibits several short but sharp oscillations, with some of these fluctuations significantly smaller in amplitude compared to others. Given these distinct patterns, a Gaussian noise assumption might not be justified, which is supported by Shapiro–Wilk’s test, yielding a $p$-value of $2.2\times 10^{-16}$. Consequently, many studies analyzing well-log data rely on a pre-cleaned dataset, wherein outliers are manually removed in advance. In general, the detection and removal of outliers are challenging tasks due to the scarcity of labelled data, stemming from the infrequent occurrence of outlier instances. Typical outlier detection methods identify data points that deviate significantly from the rest. However, proposing a universal measure of deviation applicable to all datasets and scenarios is hard or even impossible (cf. Boukerche et al., , 2020).
@@
The robustness of MUSCLE eliminates the need for a preliminary pre-cleaning step. We employ MUSCLE to estimate the median ($\beta=0.5$) of well-log signal with different choices of the tuning parameter, namely, $\alpha=0.1,0.3$ and $0.5$, and by comparison we also consider other robust or nonparametric segmentation methods, including KSD, MQS, RNSP, NOT-HT, R-FPOP and ED-PELT, see Figure 7. As shown in the top panel, the raw data has five pronounced flickering events (i.e., events on tiny temporal scales). The estimates provided by KSD, MQS, RNSP and MUSCLE with $\alpha=0.1$ or $0.3$ are very similar, and none of them detect any flickering events. Despite this, they all capture the main shape of the well-log data. We set $\alpha=0.5$ for MQS and RNSP in order to improve their detection power. By contrast, NOT-HT identifies two and R-FPOP detects three flickering events. However, R-FPOP seems to overfit the data around the last flickering (see Figure 11 in Appendix B). ED-PELT reports the highest number of change points. However, as it constantly overestimates $K$ in simulation studies, its estimation may not be trustworthy. With $\alpha=0.5$, MUSCLE successfully recovers all of the five flickering events. In general, it is challenging to verify whether these flickering events (or some of them) are genuine change points or outliers causing by mistakes in measurement. Nevertheless, MUSCLE with various choices of $\alpha$ reveals a nested sets of detected features in the well-log data. As a larger $\alpha$ corresponds a lower statistical confidence, such a nested sets of detected features shed light on how likely these features are genuine.
@@ 5.2.2 Ion channels recordings: Gramicidin D
with $\omega$ the sampling rate, and $\rho$ the convolution kernel (corresponding to the low-pass filter). The underlying signal $f$ represents the conductance profile of ion channels, which is a piecewise constant function with change points caused by gating events. The random errors $\varepsilon_{i}$ are centered, heterogeneous, and typically non-Gaussian (Venkataramanan et al., , 1998), and they are dependent due to the low-pass filter.
We showcase an extension of MUSCLE to the dependent noise in (17). To take into account the dependence structure, we need to adjust the local quantiles $q_{\alpha}$ in (4). It is generally difficult to determine the dependence of $\mathbbm{1}_{\{{\varepsilon}_{i}\leq 0\}}$, even though the dependence of noise $\varepsilon_{i}$ arises from the known low-pass filter, unless specific assumptions on the noise distribution function are made. Note however that, as a crucial building block of MUSCLE, the multiscale test statistic converges weakly to an accessible distribution in form of supremum of Gaussian random variables under mild conditions (see e.g. Frick et al., , 2014 and Dette et al., , 2020). Inspired by this, we redefine the local quantiles $q_{\alpha}$ in (4) by treating noise $\varepsilon_{i}$ as dependent Gaussian distributed, with the dependence structure derived from the convolution kernel $\rho$. We compute such local quantiles by Monte–Carlo simulations. Another challenge posed by model (17) is that the convoluted signal $\rho*f$ is no longer piecewise constant. But the influence of convolution is only locally active, as the convolution kernel $\rho$ has a compact support. As a consequence, $\rho*f$ is still constant on $s+\kappa,t)$, with $\kappa$ the kernel length (i.e. the size of the support of $\rho$), if $f$ is constant on $[s,t)$. Thus, we also modify the multiscale statistic $T_{I}$ in (3) by restricting to intervals $J$ such that $J\subseteq I$ and $J-\kappa\subseteq I$. Based on these two adjustments, we obtain an extension of MUSCLE for dependent errors, denoted by D-MUSCLE.
@@ 6 Discussion
In this paper, we introduce a novel change point segmentation method MUSCLE that is robust and meanwhile has a high detection power. It combines the ideas of multiscale testing, quantile segmentation and variational estimation. As a distinctive feature, MUSCLE controls the local errors associated with individual segments. This local error control, as opposed to the commonly studied global error control, allows improved detection power, robustness to data windowing, as well as efficient computation. Under the quantile segmentation model with serial independence, we have established the consistency and the localization error rates for MUSCLE. Remarkably, these results match (up to log factors) the minimax optimality results in Gaussian setups. Further, we obtain finite sample guarantees on its falsely detected change points in terms of FDR and OER. This provides a statistical interpretation on the unique tuning parameter $\alpha$ of MUSCLE in the sense that the expected proportion of falsely detected change points does not exceed $\alpha$. Moreover, we develop an efficient dynamic programming algorithm for MUSCLE, which leverages advanced wavelet tree data structures, modern pruning techniques, and sparse interval system proposals. We also introduce a speedup variant (i.e., MUSCLE-S) of MUSCLE, which scales linearly with sample size, particularly suitable for large scale datasets. The efficacy of MUSCLE(-S) is validated through comprehensive comparison studies using both simulated and real world data.
@@
Acknowledgments
This work was supported by DFG (German Research Foundation) CRC 1456 Mathematics of Experiment, and by the DFG under Germany’s Excellence Strategy, project EXC 2067 Multiscale Bioimaging: from Molecular Machines to Networks of Excitable Cells (MBExC). The authors thank Manuel Fink and Claudia Steinem (Institute of Organic and Biomolecular Chemistry, University of Göttingen) for providing the ion channel data, and also Axel Munk and Robin Requadt for helpful discussions.
Appendix A Proofs
We note that the statistical guarantees of MUSCLE rely solely on the fact that
$W_{i}\bigl(Z_{i},f(x_{i})\bigr)=\mathbbm{1}_{\{Z_{i}\leq f(x_{i})\}}=\mathbbm{1}_{\{\varepsilon_{i}\leq 0\}},\quad i=1,\ldots,n,$
are i.i.d. Bernoulli random variables. Therefore, the assumption of statistical independence for the observations $Z_{i}$ can be relaxed to the independence of the indicators $\mathbbm{1}_{\{Z_{i}\leq f(x_{i})\}}=\mathbbm{1}_{\{\varepsilon_{i}\leq 0\}}$.
@@ A.1 Technical results
$A_{k}(x)$ $\coloneqq\frac{x}{k}\log\left(\frac{x}{k}\cdot\frac{1}{\beta}\right)+\left(1-\frac{x}{k}\right)\log\left(\left(1-\frac{x}{k}\right)\cdot\frac{1}{1-\beta}\right)$
@@
Define a series of random times: $T_{0}\coloneqq 0$ and
$T_{s}\coloneqq\max\left\{l\,:\,\underset{T_{s-1}<i<j\leq l}{\max}B_{l-T_{s-1},i,j}-q_{\alpha}(l-T_{s-1})\leq 0\right\},$
@@
Let $S$ be in (18) and $\alpha\in(0,1)$. Then for any $s\in\mathbb{N}$, $\mathbb{P}(S\geq s)\leq\alpha^{s}.$ Moreover, we have $\mathbb{E}(S)\leq\frac{\alpha}{1-\alpha}.$
@@
$\mathbb{P}[T_{s}-T_{s-1}\geq n\mid T_{s-1}=t_{s-1},\dots,T_{1}=t_{1}]$
$=\,$ $\mathbb{P}\left[\max_{t_{s-1}<i<j\leq l}B_{l-k,i,j}\leq q_{\alpha}(l-t_{s-1}),\;\text{ for some }l\geq n+t_{s-1}\right]\geq 1-\alpha.$ (19)
Since $\mathbb{P}(S\geq s)=\mathbb{P}(T_{s}<n)=\mathbb{P}\left(\sum_{i=1}^{s}T_{s}<n\right)\leq\mathbb{P}(T_{i}-T_{i-1}<n\text{ for all }i=1,\dots,s),$ we show by introduction over $s$ that for any $n\geq 1$, $\mathbb{P}(T_{i}-T_{i-1}<n\text{ for all }i=1,\dots,s)\leq\alpha^{s}$. For $s=1$, by definitions of $q_{\alpha}(m)$ in (4) we have $\mathbb{P}(T_{1}-T_{0}<n)=\mathbb{P}(T_{1}<n)=1-\mathbb{P}(T_{1}\geq n)\leq 1-(1-\alpha)=\alpha$. Now assume that the statement holds for s1s-1, i.e., $\mathbb{P}(T_{i}-T_{i-1}<n\text{ for all }i=1,\dots,s-1)\leq\alpha^{s-1}$. Then
@@
$=\,$ $\sum_{\begin{subarray}{c}t_{i}-t_{i-1}<n\\ i=1,\dots,s-1\end{subarray}}\mathbb{P}\Bigl\{T_{s}-T_{s-1}<n,T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\dots,T_{2}-T_{1}=t_{2}-t_{1},T_{1}=t_{1}\Bigr\}$
$=\,$ $\sum_{\begin{subarray}{c}t_{i}-t_{i-1}<n\\ i=1,\dots,s-1\end{subarray}}\mathbb{P}\Bigl\{T_{s}-T_{s-1}<n\,\Big|\,T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\dots,T_{1}=t_{1}\Bigr\}$
$\mbox{}\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\times\mathbb{P}\Bigl\{T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\dots,\,T_{1}=t_{1}\Bigr\}$
$\leq\,$ $\sum_{\begin{subarray}{c}t_{i}-t_{i-1}<n\\ i=1,\dots,s-1\end{subarray}}\alpha\cdot\mathbb{P}\Bigl\{T_{s-1}-T_{s-2}=t_{s-1}-t_{s-2},\dots,T_{1}=t_{1}\Bigr\}$
@@
Let $\beta\in(0,1)$ and suppose that $Z_{1},\dots,Z_{n}$ have the common $\beta$-quantile, i.e., $\mathbb{P}(Z_{i}\leq f(x_{i}))=\beta$} with ${f}\equiv\theta$ for some $\theta\in\mathbb{R}$. Then, for any $k\geq 1$, the estimated number of change point $\widehat{K}$ by MUSCLE satisfies
$\mathbb{P}\left(\widehat{K}\geq k\right)\leq\alpha^{k}\quad\text{and}\quad\mathbb{E}\left(\widehat{K}\right)\leq\frac{\alpha}{1-\alpha}.$
@@
With loss of generality, we assume that f=0{f}=0. Suppose that there exists an estimator $\widetilde{f}=\sum_{k=0}^{\widetilde{K}}\theta_{k}I_{k}$ with $\theta_{k-1}\neq\theta_{k}$ for $k=1,\dots,K$, which has $\widetilde{K}$ change points and satisfies the multiscale side-constrains of $C_{\tilde{K}}$ in (5). Then $\widetilde{f}$ has exactly $\widetilde{K}$ false discoveries. Since the MUSCLE minimizes the number of estimated change points fulfilling the multiscale side-constrains, we have $\widehat{K}\leq\widetilde{K}$. We now construct such an estimator $\widetilde{f}$ as follows.
@@
Set $T_{0}=0$. We fit the observations $\{Z_{i}:i\geq 1\}$ using the constant value 0 up to $T_{1}$, defined as the largest index $j$ such that 0 satisfies the multiscale constraint on the interval $[0,j/n)$. We then fit the subsequent observations $\{Z_{T_{1}+i}:i\geq 1\}$ using 0 until $T_{2}$ , the largest index $j$ such that 0 satisfies the multiscale constraint on $T_{1}/n,j/n)$. Repeating this procedure until all observations are fitted yields a series of segments. For ensuring the resulting estimator has exactly $\widetilde{K}$ change points, we slightly perturb the segment values so that the multiscale constraints remain satisfied on each segment. More precisely, Let $W_{i}=\mathbbm{1}_{\{Z_{i}\leq 0\}}$ for $i=1,\dots,n$ and